OpenStudy (anonymous):

Find the rectangle of largest area that can be inscribed in a semicircle of radius R, assuming that one side of the rectangle lies on the diameter of the semicircle. NOTE: Let L denote the length of the side that lies on the diameter and H denote the height of the rectangle. Your answer will likely involve R. L= H=

OpenStudy (anonymous):

it is a calculus problem, in the chapter of optimization

OpenStudy (anonymous):



OpenStudy (anonymous):

ok

OpenStudy (anonymous):

area of that rectangle is \(A=2xy\)

OpenStudy (anonymous):

since \(x^2+y^2=r^2\) you have \[y=\sqrt{r^2-x^2}\]

OpenStudy (anonymous):

that makes \[A(x)=2x\sqrt{r^2-x^2}\] etc

OpenStudy (anonymous):

@satellite73 so what would be next?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

take the derivative, find the critical points etc

OpenStudy (anonymous):

ok im taking the derivative right nw

OpenStudy (anonymous):

(2 (r^2 - 2 x^2))/Sqrt[r^2 - x^2]

OpenStudy (anonymous):

you need the product rule for it maybe... let me check

OpenStudy (anonymous):

yeah that looks good

OpenStudy (anonymous):

when i make it equal to zero to then find the critical points, i get r/sqrt(2) and -r/sqrt(2)

OpenStudy (anonymous):

set \[r^2-2x^2=0\] solve for \(x\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yeah that looks good also

OpenStudy (anonymous):

so whats next?

OpenStudy (anonymous):

after the critical points

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

well i guess i don't mean nothing lets go back to the picture you have found \(x=\frac{r}{\sqrt2}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

you are asked for \(H\) which is evidently \[\frac{2r}{\sqrt2}=\sqrt2r\]

OpenStudy (anonymous):

and also \(L\) which we called \(y\) and you can solve for since you know \[x^2+y^2=r^2\]

OpenStudy (anonymous):

so H is? and L?

OpenStudy (anonymous):

i guess i confused you i said what \(H\) is above

OpenStudy (anonymous):

H = sqrt(2)r ?

OpenStudy (anonymous):

I answered that for H and i got it wrong

OpenStudy (anonymous):

oh because i am an idiot and don't know how to read L is the length, H is the height (doh)

OpenStudy (anonymous):

that makes \(L=\sqrt2 r\)

OpenStudy (anonymous):

and H?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

I need help finding H, @satellite73

OpenStudy (anonymous):

@dan815 help me

OpenStudy (anonymous):

@ganeshie8